OpenStudy (anonymous):

what is the derivative of f(x)= 3x^2+3 at x=2?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

12

OpenStudy (anonymous):

d/dx of (3(x^2)+3) = 6x so at x=2, derivative 6*2=12

OpenStudy (anonymous):

y=3x^2+3 dy/dx= 3 d/dx x^2+ d/dx 3 dy/dx = 3(2x)+0 dy/dx = 6x at x=2 6*2=12

OpenStudy (anonymous):

yeah 12 is correct like shristi1 showed

OpenStudy (anonymous):

thanks ! : )

OpenStudy (anonymous):

12 as dy/dx gives 6x which then produces 12

OpenStudy (anonymous):

y=3x^2+3 dy/dx= d/dx(3x^2+3) dy/dx = 3(2x)+0 dy/dx = 6x at x=2 6(2)=12 Answer = 12